Neutron stars are some of the densest objects in the universe. They form during supernova ex plosions and are very small compared to other astronomical objects. You can assume 5 * 10 ^ 17 kg/m³ as the average density of a neutron star. The Earth has a total mass of about 5.97*10^ 24 kg


What would be the diameter of the Earth if it had the density of a neutron star?


1
Expert's answer
2022-03-01T11:12:26-0500

Let us determine the volume of neutron star: "V=\frac43 \pi R^3." We should obtain an estimate for R.

The density "\rho = \dfrac{M}{V}," so "V=\dfrac{M} {\rho}."


The volume will be "V=\dfrac{5.97\cdot10^{24}}{5\cdot10^{17}} = 1.19\cdot10^7\,\mathrm{m^3}."


The radius will be "R= \sqrt[3]{\dfrac{V}{\frac43\pi}} = 142\,\mathrm{m}."

The diameter will be 284 m.